# Equilibrium Policy Generalization: A Reinforcement Learning Framework for Cross-Graph Zero-Shot Generalization in Pursuit-Evasion Games

### v1

@@ 1 Introduction
Real-world environments are variable, and the dynamics for real-world games can be time-evolving. As an important class of real-world games, pursuit-evasion games (PEGs) can model a variety of real-world problems (e.g., in robotics and security domains [28; 29; 7]). Typically with a team of pursuers and an adversarial evader, PEGs can be formulated with complex graph structures, where nodes and edges can be temporarily removed or added as the game proceeds. An intelligent game agent should be robust to such changes. In other words, an ideal pursuer or evader policy should be robust to different map or graph inputs. Classical differential game [22; 35] and dynamic programming [30; 13] methods, while accurate, require recomputation under potential changes to game dynamics. In view of this gap, recent works on graph-based PEGs [16; 17; 32; 33] employ deep reinforcement learning (RL) to obtain a more flexible policy. Among the state-of-the-art methods, Grasper (see [16]) can pretrain a generalized multi-agent pursuer policy robust to different initial conditions (e.g., exit positions). Given an actual game situation, subsequent fine-tuning through game-theoretic approaches like PSRO [15] can be used to improve the policy without recomputing from scratch.
However, the state-of-the-art methods still exhibit two practical limitations. First, real-time applicability remains questionable due to the requirement of computationally intensive policy fine-tuning. Recent work [37] points out that existing methods may struggle to adapt to rapid changes in urban settings (e.g., traffic jams, emergencies, and unexpected social events). To make it worse, when the graph structure significantly changes, the pretrained policy will lose the guarantee to be a good starting point for subsequent tuning. Second, existing training processes heavily rely on certain behavior patterns of the opponent policy, which impairs the robustness of the methods. For example, Grasper requires the evader to consistently follow the shortest path to the chosen exits. As a result, the method is not applicable to no-exit PEGs and not general enough for both sides of the players. Besides, such simplifications can make the RL policy more exploitable by an intelligent opponent.
@@
  • We propose Equilibrium Policy Generalization (EPG), a novel reinforcement learning framework that enables zero-shot generalization in Markov PEGs across different graph structures. We are the first to show that reinforcement learning on a corpus of graphs with equilibrium policies serving as adversaries (and also guidance) can lead to a generalized pursuer (or evader) policy with robust zero-shot performance in unseen real-world graphs.
  • For no-exit graphs, we present a dynamic programming (DP) algorithm to efficiently construct an oracle that generates the single-graph equilibrium policies in Markov PEGs. We prove that the proposed algorithm can compute pure-strategy Nash equilibrium under a near-optimal $\tilde{\mathcal{O}}(\left|S\right|)$ time complexity. We further extend this algorithm with a grouping mechanism to facilitate scalability with respect to pursuer number.
  • We design a decentralized network architecture with a novel shortest path distance feature as state input to represent cross-graph multi-agent PEG policies and verify the effectiveness of the distance feature for EPG training. Source code is in the Supplementary Material (with DP implementation and test files) and https://github.com/Cahemgco/EPG_code (simplified).
  • For graphs with exits, we provide a heuristic approach based on bipartite graph matching as an effective substitute for an exact equilibrium oracle. Under such an approximation, we verify that the EPG framework still guarantees a strong zero-shot performance that matches or outperforms the fine-tuned results from the state-of-the-art methods like Grasper [16].
@@ 2.1 Markov Game and Nash Equilibrium
Nash equilibrium. A Nash equilibrium (NE) in a game is a joint policy where each individual player cannot benefit from unilaterally deviating from his/her own policy. Specifically, in a two-player zero-sum MG, an NE $(\mu^{*},\nu^{*})$ satisfies ${{V}^{\mu,{{\nu}^{*}}}}\leq{{V}^{{{\mu}^{*}},{{\nu}^{*}}}}\leq{{V}^{{{\mu}^{*}},\nu}}$ for any $\mu$ and $\nu$ at all states. As is well known, every MG with finite states and actions has at least one NE, and all NEs in a two-player zero-sum MG share the same Nash value ${{V}^{*}}(s)={{V}^{{{\mu}^{*}},{{\nu}^{*}}}}(s)={{\max}_{\mu}}{{\min}_{\nu}}{{V}^{\mu,\nu}}(s)={{\min}_{\nu}}{{\max}_{\mu}}{{V}^{\mu,\nu}}(s)$ [25]. In two-player zero-sum games, Nash equilibrium can be viewed as the optimal joint policy since both players cannot be exploited by their worst-case opponents. Besides, if $(\mu_{1},\nu_{1})$ and $(\mu_{2},\nu_{2})$ are both NEs, then $(\mu_{1},\nu_{2})$ and $(\mu_{2},\nu_{1})$ are NEs as well [24]. To guarantee the optimality, it suffices to find an equilibrium policy for each player.
@@ 3.1 Reinforcement Learning with Equilibrium Adversary
SAC [11; 5] is a well-known off-policy reinforcement learning framework that maximizes expected return regularized under a policy entropy term. Specifically, the value function in discrete-action SAC [5] is defined as $V(s)={{\mathbb{E}}_{a\sim\pi(s)}}\left[Q(s,a)-\alpha\log\pi(s,a)\right]$. The losses of the value network $Q_{\phi}$ and the policy network $\pi_{\theta}$ are computed as ${{J}_{Q}}(\phi)={{\mathbb{E}}_{s,a}}\left[\frac{1}{2}{{\left({{Q}_{\phi}}(s,a)-\left(r+\gamma{{\mathbb{E}}_{{{s}^{\prime}}}}\left[V({{s}^{\prime}})\right]\right)\right)}^{2}}\right]$ and ${{J}_{\pi}}(\theta)={{\mathbb{E}}_{s,a\sim{{\pi}_{\theta}}(s)}}\left[\alpha\log{{\pi}_{\theta}}(s,a)-Q(s,a)\right]$, respectively. Besides, the temperature $\alpha$ under target entropy $\overline{H}$ is adaptively updated under loss $J(\alpha)={{\mathbb{E}}_{s,a}}\left[-\alpha\left(\log\pi(s,a)+\overline{H}\right)\right]$. Here, we use SAC as the foundation algorithm to introduce the RL pipeline in Equilibrium Policy Generalization.
@@ 3.2.1 Dynamic Programming for No-Exit Markov PEGs
Determining whether the pursuers can always capture the evader in a no-exit PEG is proved to be an EXPTIME-complete problem (see [10]). This result suggests that an accurate equilibrium oracle could incur a time complexity of $\Omega(|S|)$, where $|S|$ is the number of all game states, which is exponential in the agent number $m+1$. For sequential PEGs under no-exit graphs, [30] proposes an efficient method to compute the minimum steps for the pursuers to capture the evader. The idea is to iteratively expand the set of states where the pursuit is guaranteed to be successful under the optimal strategy. During the $\mathcal{O}(|S|)$ state expansion, the equilibrium policy is also generated. For Markov games, however, the method is not applicable since the moves for both players are simultaneous.
On the other hand, the classical marking algorithm (see [7]) can have a worst-case time complexity of $\mathcal{O}(|S|^{2})$ in practical Markov PEGs. Even when the pursuer number is small (e.g., two or three), the algorithm cannot scale with the number of nodes in the graph, which can be large in the real world. In view of this gap, we first show how to apply the idea of state expansion to simultaneous games and provide a more efficient dynamic programming (DP) algorithm for Markov PEGs.
@@ 17         Push $(n_{p},n_{e})$ into $\mathcal{Q}$
The detailed proof of Theorem 1 is based on mathematical induction and reserved in Appendix A.1. Actually, Algorithm 1 efficiently solves the Bellman minimax equation (see [36]) under the existence of a pure-strategy Nash equilibrium. Note that the update condition $D(s_{p},s_{e})=\infty$ guarantees that every state $(s_{p},s_{e})$ is pushed into and popped from $\mathcal{Q}$ at most once. Therefore, the time complexity of Algorithm 1 is a near-optimal $\tilde{\mathcal{O}}(\left|S\right|)$, where $\tilde{\mathcal{O}}$ hides the logarithm factors required for picking each $(n_{p},n_{e})$ through preserving data structures like balanced trees. Also note that the computation of $\mu(s)$ and $\nu(s)$ can be reserved online and thus does not affect the overall time complexity.
Algorithm 1 suggests that the Nash value admits an efficient estimation when the PEG only admits a one-sided termination function. Compared to direct value iteration (see [13]), the DP algorithm can exactly compute the equilibrium policy within finite iterations. The existence of pure-strategy Nash equilibrium implies that a successful pursuit is guaranteed for all $s\in S$ (i.e., $D(s)<\infty$). For example, the optimal strategy in any tree-form $G$ is a pure-strategy Nash equilibrium if we regard adjacency as the condition of a successful pursuit. Even when this assumption does not globally hold in a PEG, Algorithm 1 still induces a near-optimal pursuit strategy for the states with finite $D(\cdot)$.
@@ 3.3 Cross-Graph Representation of PEG Policy
With the above-mentioned considerations, we present a sequence model with an attention-based architecture (see Figure 2) to represent the cross-graph joint policy. For a team of pursuers with $m$ agents, the sequence model queries the policy network $m$ times under a fixed adjacent matrix $M\in\{0,1\}^{n\times n}$ ($n=\left|\mathcal{V}\right|$) for the current graph. The input is composed of a state feature $s_{f}$ that describes the current global state and the information of node index $c$ for the current acting agent. To uniformly capture the state information across different graphs, we use the shortest path distances to the $m+1$ agents (including the evader) and the exits (if existing) as the initial feature of each node $v\in\mathcal{V}$. The state feature $s_{f}$ is composed of the normalized features of all $n$ nodes. The shortest path is an important and well-examined concept in graph theory [2], and the distance between any two nodes can be preprocessed using the $\mathcal{O}(n^{3})$ Floyd algorithm. In Appendix A.2, we further prove that the input $(s_{f},c)$ is dense enough to identify the current global state in sequential decision-making.
@@
Given the state feature input $s_{f}$, we borrow the ideas from the existing works on robot exploration [3; 18] to construct a policy network (on the right of Figure 2). We first embed $s_{f}$ into $\mathbb{R}^{d\times n}$ and send it into an encoder composed of multiple self-attention [27] layers, where $d$ is the embedding dimension. Each layer takes the output $h$ of the last layer as the input and outputs $h^{\prime}$ using a masked attention: ${{q}_{i}}={{W}_{Q}}{{h}_{i}},{{k}_{i}}={{W}_{K}}{{h}_{i}},\ {{v}_{i}}={{W}_{V}}{{h}_{i}},{{u}_{ij}}=\frac{q_{i}^{T}{{k}_{j}}}{\sqrt{d}},{{w}_{ij}}=\frac{{{e}^{{{u}_{ij}}}}}{\sum\nolimits_{t=1}^{n}{{{e}^{{{u}_{it}}}}}},h_{i}^{\prime}=\sum\limits_{j=1}^{n}{\min\left\{{{w}_{ij}},{{M}_{ij}}\right\}{{v}_{j}}}$, where $W_{Q},W_{K},W_{V}\in\mathbb{R}^{d\times d}$ are the weights to be learned. Theoretically, it requires $\textbf{Diameter}(G)$ layers to globally broadcast the local information of a node since $M$ is the adjacent matrix. However, in contrast to using a low-level node feature with indicators (see Grasper [16]), using a high-level distance feature allows the node itself to encode long-term information. With the shortest path distance feature, we can use a fixed number of attention layers ($6$ in practice) for information transmission.
Denote by $\hat{h}$ the output of the encoder, and recall that $c$ is the node index corresponding to the current acting agent. We further use a decoder without masking to gather global information. Specifically, the decoder uses $\hat{h}_{c}$ to query in the output features $\hat{h}$ of all nodes, with the keys equal to the values: $q={{W}_{Q}}{{{\hat{h}}}_{c}},{{k}_{i}}={{W}_{K}}{{{\hat{h}}}_{i}},{{v}_{i}}={{W}_{V}}{{{\hat{h}}}_{i}},{{u}_{j}}=\frac{{{q}^{T}}{{k}_{j}}}{\sqrt{d}},{{w}_{j}}=\frac{{{e}^{{{u}_{j}}}}}{\sum\nolimits_{t=1}^{n}{{{e}^{{{u}_{t}}}}}},{\tilde{h}_{c}}=\sum\nolimits_{j=1}^{n}{{{w}_{j}}{{v}_{j}}}$. The decoder output $\tilde{h}_{c}$ is further concatenated with $\hat{h}_{c}$ and projected into $\mathbb{R}^{d}$. Then, it is used as a query for a pointer network [31], which takes the features of the neighbor nodes $\hat{h}_{ne}$ for the current agent as the keys and values. The pointer network directly outputs the attention vector $w$ as the current policy (i.e., $\pi(a|s)=w_{a}$) since the number of the neighbors aligns with the number of the valid actions.

### v2

@@ 1 Introduction
Real-world environments are variable, and the dynamics for real-world games can be time-evolving. As an important class of real-world games, pursuit-evasion games (PEGs) can model a variety of real-world problems (e.g., in robotics and security domains [vidal2001pursuit; vidal2002probabilistic; chung2011search]). Typically with a team of pursuers and an adversarial evader, PEGs can be formulated with complex graph structures, where nodes and edges can be temporarily removed or added as the game proceeds. An intelligent game agent should be robust to such changes. In other words, an ideal pursuer or evader policy should be robust to different map or graph inputs. Classical differential game [margellos2011hamilton; zhou2012general] and dynamic programming [vieira2008optimal; horak2017dynamic] methods, while accurate, require recomputation under potential changes to game dynamics. In view of this gap, recent works on graph-based PEGs [li2024grasper; li2023solving; xue2022nsgzero; xue2021solving] employ deep reinforcement learning (RL) to obtain a more flexible policy. Among the state-of-the-art methods, Grasper (see [li2024grasper]) can pretrain a generalized multi-agent pursuer policy robust to different initial conditions (e.g., exit positions). Given an actual game situation, subsequent fine-tuning through game-theoretic approaches like PSRO [lanctot2017unified] can be used to improve the policy without recomputing from scratch.
However, the state-of-the-art methods still exhibit two practical limitations. First, real-time applicability remains questionable due to the requirement of computationally intensive policy fine-tuning. Recent work [zhuang2025solving] points out that existing methods may struggle to adapt to rapid changes in urban settings (e.g., traffic jams, emergencies, and unexpected social events). To make it worse, when the graph structure significantly changes, the pretrained policy will lose the guarantee to be a good starting point for subsequent tuning. Second, existing training processes heavily rely on certain behavior patterns of the opponent policy, which impairs the robustness of the methods. For example, Grasper requires the evader to consistently follow the shortest path to the chosen exits. As a result, the method is not applicable to no-exit PEGs and not general enough for both sides of the players. Besides, such simplifications can make the RL policy more exploitable by an intelligent opponent.
@@
  • We propose Equilibrium Policy Generalization (EPG), a novel reinforcement learning framework that enables zero-shot generalization in Markov PEGs across different graph structures. For the first time, we show that reinforcement learning on a corpus of graphs with equilibrium policies serving as adversaries (and also guidance) can lead to a generalized pursuer (or evader) policy with robust zero-shot performance in unseen real-world graphs.
  • For no-exit graphs, we present a dynamic programming (DP) algorithm to efficiently construct an oracle that generates the single-graph equilibrium policies in Markov PEGs. We prove that the proposed algorithm can compute pure-strategy Nash equilibrium under a near-optimal $\tilde{\mathcal{O}}(\left|S\right|)$ time complexity. We further extend this algorithm with a grouping mechanism for scalability with respect to pursuer number.
  • We design a decentralized network architecture with a novel shortest path distance feature as state inputs to represent cross-graph multi-agent PEG policies. Through ablation studies, we verify that the distance feature, along with our reinforcement learning loss, plays an important role in cross-graph EPG training. Source code can be found in our supplementary material (with DP implementation and test files) and at https://github.com/Cahemgco/EPG_code.
  • For graphs with exits, we provide a heuristic approach based on bipartite graph matching as an effective substitute for an exact equilibrium oracle. Directly utilizing this heuristic and approximate equilibrium oracle, we verify that the EPG framework still guarantees a strong zero-shot performance that matches or outperforms the fine-tuned results from the state-of-the-art methods like Grasper [li2024grasper].
@@ 2.1 Markov Game and Nash Equilibrium
Nash equilibrium. A Nash equilibrium (NE) in a game is a joint policy where each individual player cannot benefit from unilaterally deviating from his/her own policy. Specifically, in a two-player zero-sum MG, an NE $(\mu^{*},\nu^{*})$ satisfies ${{V}^{\mu,{{\nu}^{*}}}}\leq{{V}^{{{\mu}^{*}},{{\nu}^{*}}}}\leq{{V}^{{{\mu}^{*}},\nu}}$ for any $\mu$ and $\nu$ at all states. As is well known, every MG with finite states and actions has at least one NE, and all NEs in a two-player zero-sum MG share the same Nash value ${{V}^{*}}(s)={{V}^{{{\mu}^{*}},{{\nu}^{*}}}}(s)={{\max}_{\mu}}{{\min}_{\nu}}{{V}^{\mu,\nu}}(s)={{\min}_{\nu}}{{\max}_{\mu}}{{V}^{\mu,\nu}}(s)$ [shapley1953stochastic]. In two-player zero-sum games, Nash equilibrium can be viewed as the optimal joint policy since both players cannot be exploited by their worst-case opponents. Besides, if $(\mu_{1},\nu_{1})$ and $(\mu_{2},\nu_{2})$ are both NEs, then $(\mu_{1},\nu_{2})$ and $(\mu_{2},\nu_{1})$ are NEs as well [roughgarden2016twenty]. To guarantee the optimality, it suffices to find an equilibrium policy for each player.
@@ 3.1 Reinforcement Learning with Equilibrium Adversary
SAC [haarnoja2018soft; christodoulou2019soft] is a well-known off-policy reinforcement learning framework that maximizes expected return regularized under a policy entropy term. Specifically, the value function in discrete-action SAC [christodoulou2019soft] is defined as $V(s)={{\mathbb{E}}_{a\sim\pi(s)}}\left[Q(s,a)-\alpha\log\pi(s,a)\right]$. The losses of the value network $Q_{\phi}$ and the policy network $\pi_{\theta}$ are computed as ${{J}_{Q}}(\phi)={{\mathbb{E}}_{s,a}}\left[\frac{1}{2}{{\left({{Q}_{\phi}}(s,a)-\left(r+\gamma{{\mathbb{E}}_{{{s}^{\prime}}}}\left[V({{s}^{\prime}})\right]\right)\right)}^{2}}\right]$ and ${{J}_{\pi}}(\theta)={{\mathbb{E}}_{s,a\sim{{\pi}_{\theta}}(s)}}\left[\alpha\log{{\pi}_{\theta}}(s,a)-Q(s,a)\right]$, respectively. Besides, the temperature $\alpha$ under target entropy $\overline{H}$ is adaptively updated under loss $J(\alpha)={{\mathbb{E}}_{s,a}}\left[-\alpha\left(\log\pi(s,a)+\overline{H}\right)\right]$. Here, we use SAC as the foundation algorithm to introduce the RL pipeline in Equilibrium Policy Generalization.
@@ 3.2.1 Dynamic Programming for No-Exit Markov PEGs
Determining whether the pursuers can always capture the evader in a no-exit PEG is proved to be an EXPTIME-complete problem (see [goldstein1995complexity]). This result suggests that an accurate equilibrium oracle could incur a time complexity of $\Omega(|S|)$, where $|S|$ is the number of all game states, which is exponential in the agent number $m+1$. For sequential PEGs under no-exit graphs, [vieira2008optimal] proposes an efficient method to compute the minimum steps for the pursuers to capture the evader. The idea is to iteratively expand the set of states where the pursuit is guaranteed to be successful under the optimal strategy. During the $\mathcal{O}(|S|)$ state expansion, the equilibrium policy is also generated. For Markov games, however, the method is not applicable since the moves for both players are simultaneous.
On the other hand, the classical marking algorithm (see [chung2011search]) can have a worst-case time complexity of $\mathcal{O}(|S|^{2})$ in practical Markov PEGs. Even when the pursuer number is small (e.g., two or three), the algorithm cannot scale with the number of nodes in the graph, which can be large in the real world. In view of this gap, we first show how to apply the idea of state expansion to simultaneous games and provide a more efficient dynamic programming (DP) algorithm for Markov PEGs.
@@ 17         Push $(n_{p},n_{e})$ into $\mathcal{Q}$
The detailed proof of Theorem 1 is based on mathematical induction and reserved in Appendix A.1. Actually, Algorithm 1 efficiently solves the Bellman minimax equation (see [zhu2020online]) under the existence of a pure-strategy Nash equilibrium. Note that the update condition $D(s_{p},s_{e})=\infty$ guarantees that every state $(s_{p},s_{e})$ is pushed into and popped from $\mathcal{Q}$ at most once. Therefore, the time complexity of Algorithm 1 is a near-optimal $\tilde{\mathcal{O}}(\left|S\right|)$, where $\tilde{\mathcal{O}}$ hides the logarithm factors required for picking each $(n_{p},n_{e})$ through preserving data structures like balanced trees. Also note that the computation of $\mu(s)$ and $\nu(s)$ can be reserved online and thus does not affect the overall time complexity.
Algorithm 1 suggests that the Nash value admits an efficient estimation when the PEG only admits a one-sided termination function. Compared to direct value iteration (see [horak2017dynamic]), the DP algorithm can exactly compute the equilibrium policy within finite iterations. The existence of pure-strategy Nash equilibrium implies that a successful pursuit is guaranteed for all $s\in S$ (i.e., $D(s)<\infty$). For example, the optimal strategy in any tree-form $G$ is a pure-strategy Nash equilibrium if we regard adjacency as the condition of a successful pursuit. Even when this assumption does not globally hold in a PEG, Algorithm 1 still induces a near-optimal pursuit strategy for the states with finite $D(\cdot)$.
@@ 3.3 Cross-Graph Representation of PEG Policy
With the above-mentioned considerations, we present a sequence model with an attention-based architecture (see Figure 2) to represent the cross-graph joint policy. For a team of pursuers with $m$ agents, the sequence model queries the policy network $m$ times under a fixed adjacent matrix $M\in\{0,1\}^{n\times n}$ ($n=\left|\mathcal{V}\right|$) for the current graph. The input is composed of a state feature $s_{f}$ that describes the current global state and the information of node index $c$ for the current acting agent. To uniformly capture the state information across different graphs, we use the shortest path distances to the $m+1$ agents (including the evader) and the exits (if existing) as the initial feature of each node $v\in\mathcal{V}$. The state feature $s_{f}$ is composed of the normalized features of all $n$ nodes. The shortest path is an important and well-examined concept in graph theory [bondy2008graph], and the distance between any two nodes can be preprocessed using the $\mathcal{O}(n^{3})$ Floyd algorithm. In Appendix A.2, we further prove that the input $(s_{f},c)$ is dense enough to identify the current global state in sequential decision-making.
@@
Given the state feature input $s_{f}$, we borrow the ideas from the existing works on robot exploration [cao2023ariadne; liu2025two] to construct a policy network (on the right of Figure 2). We first embed $s_{f}$ into $\mathbb{R}^{d\times n}$ and send it into an encoder composed of multiple self-attention [vaswani2017attention] layers, where $d$ is the embedding dimension. Each layer takes the output $h$ of the last layer as the input and outputs $h^{\prime}$ using a masked attention: ${{q}_{i}}={{W}_{Q}}{{h}_{i}},{{k}_{i}}={{W}_{K}}{{h}_{i}},\ {{v}_{i}}={{W}_{V}}{{h}_{i}},{{u}_{ij}}=\frac{q_{i}^{T}{{k}_{j}}}{\sqrt{d}},{{w}_{ij}}=\frac{{{e}^{{{u}_{ij}}}}}{\sum\nolimits_{t=1}^{n}{{{e}^{{{u}_{it}}}}}},h_{i}^{\prime}=\sum\limits_{j=1}^{n}{\min\left\{{{w}_{ij}},{{M}_{ij}}\right\}{{v}_{j}}}$, where $W_{Q},W_{K},W_{V}\in\mathbb{R}^{d\times d}$ are the weights to be learned. Theoretically, it requires $\textbf{Diameter}(G)$ layers to globally broadcast the local information of a node since $M$ is the adjacent matrix. However, in contrast to using a low-level node feature with indicators (see Grasper [li2024grasper]), using a high-level distance feature allows the node itself to encode long-term information. With the shortest path distance feature, we can use a fixed number of attention layers ($6$ in practice) for information transmission.
Denote by $\hat{h}$ the output of the encoder, and recall that $c$ is the node index corresponding to the current acting agent. We further use a decoder without masking to gather global information. Specifically, the decoder uses $\hat{h}_{c}$ to query in the output features $\hat{h}$ of all nodes, with the keys equal to the values: $q={{W}_{Q}}{{{\hat{h}}}_{c}},{{k}_{i}}={{W}_{K}}{{{\hat{h}}}_{i}},{{v}_{i}}={{W}_{V}}{{{\hat{h}}}_{i}},{{u}_{j}}=\frac{{{q}^{T}}{{k}_{j}}}{\sqrt{d}},{{w}_{j}}=\frac{{{e}^{{{u}_{j}}}}}{\sum\nolimits_{t=1}^{n}{{{e}^{{{u}_{t}}}}}},{\tilde{h}_{c}}=\sum\nolimits_{j=1}^{n}{{{w}_{j}}{{v}_{j}}}$. The decoder output $\tilde{h}_{c}$ is further concatenated with $\hat{h}_{c}$ and projected into $\mathbb{R}^{d}$. Then, it is used as a query for a pointer network [vinyals2015pointer], which takes the features of the neighbor nodes $\hat{h}_{ne}$ for the current agent as the keys and values. The pointer network directly outputs the attention vector $w$ as the current policy (i.e., $\pi(a|s)=w_{a}$) since the number of the neighbors aligns with the number of the valid actions.